OpenStudy (anonymous):

Tan-1(1+x/1-x)^1/2find dy/dx

OpenStudy (anonymous):

arctan sqrt[(1+x)/(1-x)] dy/dx = {sqrt[(1+x)/(1-x)]}'/{1+[sqrt[(1+x)/(1-x)]]^2} dy/dx = [(1+x)/(1-x)]'/2{sqrt[(1+x)/(1-x)]}*{1+(1+x)/(1-x)} [(1+x)/(1-x)]' = (1-x-1-x)/(1-x)^2=-2x/(1-x)^2 dy/dx = -2x(1-x)/4*(1-x)^2*{sqrt[(1+x)/(1-x)]} dy/dx = -x/2*(1-x)*{sqrt[(1+x)/(1-x)]}